OpenStudy (anonymous):

Rewrite the following equation in the center-radius form of the equation of a circle. x2 + 10x = -y2 + 12y + 3?

OpenStudy (anonymous):

\[x^2 + 10x = -y^2 + 12y + 3\]

OpenStudy (anonymous):

start with \[x^2+10x+y^2-12y=3\] then complete the square twice, once for the x terms and once for the y terms do you know how to complete the square?